Crenshaw Enterprises has gathered projected cash flows for two projects.
Year Project I Project J
0 –$255,000 –$255,000
1 114,500 88,000
2 104,000 99,500
3 88,000 101,500
4 77,000 108,500

a.

At what interest rate would the company be indifferent between the two projects? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)

b. Which project is better if the required return is above this interest rate?
Answer #1

a) We need to use excel's goalseek function or a financial calculator to find the discount rate at which the NPV of the two projects is the same and then the investor will be indifferent between the two:

Using excel we calculated the discount rate as 15.75% rounded to 2 decimal places

Project I

Year CF Discount Factor Discounted CF
0 $ -2,55,000.00 1/(1+0.157502319529157)^0= 1 1*-255000= $ -2,55,000.00
1 $ 1,14,500.00 1/(1+0.157502319529157)^1= 0.863929154 0.863929154290399*114500= $       98,919.89
2 $ 1,04,000.00 1/(1+0.157502319529157)^2= 0.746373584 0.746373583632924*104000= $       77,622.85
3 $      88,000.00 1/(1+0.157502319529157)^3= 0.644813899 0.644813898892686*88000= $       56,743.62
4 $      77,000.00 1/(1+0.157502319529157)^4= 0.557073526 0.557073526345053*77000= $       42,894.66
NPV = Sum of all Discounted CF $       21,181.03

Project J

Year CF Discount Factor Discounted CF
0 $ -2,55,000.00 1/(1+0.157502319529157)^0= 1 1*-255000=      -2,55,000.00
1 $      88,000.00 1/(1+0.157502319529157)^1= 0.863929154 0.863929154290399*88000=           76,025.77
2 $      99,500.00 1/(1+0.157502319529157)^2= 0.746373584 0.746373583632924*99500=           74,264.17
3 $ 1,01,500.00 1/(1+0.157502319529157)^3= 0.644813899 0.644813898892686*101500=           65,448.61
4 $ 1,08,500.00 1/(1+0.157502319529157)^4= 0.557073526 0.557073526345053*108500=           60,442.48
NPV = Sum of all Discounted CF           21,181.03

b) Using this knowledge, we take the discount rate as 16% and recalculate the NPV: and it comes to:

Project I:

Year CF Discount Factor Discounted CF
0 $ -2,55,000.00 1/(1+0.16)^0= 1 1*-255000= $ -2,55,000.00
1 $ 1,14,500.00 1/(1+0.16)^1= 0.862068966 0.862068965517241*114500= $       98,706.90
2 $ 1,04,000.00 1/(1+0.16)^2= 0.743162901 0.743162901307967*104000= $       77,288.94
3 $      88,000.00 1/(1+0.16)^3= 0.640657674 0.640657673541351*88000= $       56,377.88
4 $      77,000.00 1/(1+0.16)^4= 0.552291098 0.552291097880475*77000= $       42,526.41
NPV = Sum of all Discounted CF $       19,900.13

Project J:

Year CF Discount Factor Discounted CF
0 $ -2,55,000.00 1/(1+0.16)^0= 1 1*-255000=      -2,55,000.00
1 $      88,000.00 1/(1+0.16)^1= 0.862068966 0.862068965517241*88000=           75,862.07
2 $      99,500.00 1/(1+0.16)^2= 0.743162901 0.743162901307967*99500=           73,944.71
3 $ 1,01,500.00 1/(1+0.16)^3= 0.640657674 0.640657673541351*101500=           65,026.75
4 $ 1,08,500.00 1/(1+0.16)^4= 0.552291098 0.552291097880475*108500=           59,923.58
NPV = Sum of all Discounted CF           19,757.12

So we see that Project I is better and this is the case for other higher rates also:

Discount Rate Project I Project J
16.00% $   19,900.13 $   19,757.12
17.00% $   14,872.32 $   14,174.42
18.00% $   10,000.34 $     8,774.75
19.00% $     5,277.71 $     3,550.24
20.00% $         698.30 $   -1,506.56